About ​8% of the population of a large country is . If two people are randomly​ selected, what is the probability both are ​? What is the probability at least one is ​?
1
Expert's answer
2020-11-02T20:46:16-0500

If two people are randomly​ selected, then the probability that both have that feature is:

"P(2) = 0.08^2 = 0.0064."

​The probability that at least one of them has that feature is:

"P(1 or 2) = 1 - (1 - 0.08)^2 = 0.1536."